# Enhancing Video Large Language Models with Structured Multi-Video Collaborative Reasoning (early version)

### v1

@@ 3.4 Structured Multi-Video Prompt
Upon obtaining the fused multi-video graph tokens, we need to integrate the graph, video, and text tokens together to create an LLM-friendly prompt. Therefore, we propose the structured multi-video prompt, as is shown in Figure 4. Our prompt originates from previous video language models’ prompt design [28]. For target video, we maintain the video tokens <<VIDEO_TOKENS>> of the target video to preserve fine-grained and detailed visual information. We also append its graph-structured data <<GRAPH_TOKENS>> to indicate the key object and spatio-temporal relationships. For $N$ related videos, we only maintain the concise and data-efficient graph-structured data <<GRAPH_TOKENS>>. In this context, we further indicate the relationships between the target video and related videos and how the LLMs can utilize these related multi-video structured data. Moreover, we introduce Chain-of-Thought (CoT) [49] reasoning through the fixed prompt “Think through the process step by step”. This approach helps the model to sequentially reason through structured information, improving its ability to comprehend and process the graph-based features. By structuring the prompt in this manner, we enable the video language model to leverage multi-video information in an effective way, thereby enhancing the model’s ability to reason about and answer queries related to the video content.
@@ 4.1 Experiment Setup
Dataset descriptions. At the training stage, we finetune our model on a subset of VideoInstruct-100K [36] dataset with 10K video-text instructions together with the structured video graph. For evaluation, we test our method on three Video-QA datasets, including two public open-ended video question-answering datasets: ActivityNet-QA [59] and MSRVTT-QA [54]. Besides, we also establish a high-quality domain-related Video-QA dataset named InternVid-QA. We meticulously selected 631 videos across 41 major domains from the large-scale video dataset InternVid [47]. We illustrated the statistics and some video examples in Figure 5. InternVid-QA’s videos cover various domains of life (see Figure 5(a)) and are annotated with high-quality, domain-relevant question-answer pairs (see Figure 5(b)). InternVid-QA’s setting aligns more closely with the objectives of our multi-video collaborative understanding. InternVid-QA will be public available for research use.
Implementation details. Our proposed structured multi-video collaboration framework is adaptive to general video language models. To conduct experiments, we uses Vicuna-7B v1.5 [7] as the language model and LanguageBind [64] based on ViT-L/14 [10] as its video encoder, following Video-LLaVA [28]. For the Graph Fusion Module, the output node-level feature’s dimension is 1024. Since Video-LLaVA has pre-trained the projection layer to align the OpenCLIP’s visual features with LLM’s text tokens, we just pass the 1024-dim node feature to the projector and get the 4096-dim graph tokens. We initialize our video language model with Video-LLaVA 7B’s model weights and empower the model with the capability to learn from multi-video graphs through finetuning. In the fine-tuning process, the vision and text encoder are frozen. For the remaining network, we use the AdamW optimizer with a cosine learning rate schedule, whose initial learning rate is $2\times 10^{-5}$ and the warmup ratio is 0.03. We train our model on 8 A6000 GPUs with a total batch size of 32. For each video, we uniformly sample 8 frames and resize each frame to $224\times 224$. Each video is accompanied by 5 related videos during training and experiments are conducted on InternVid-QA without special comments.
Figure 5: The InternVid-QA dataset. (a) Statistic results of video number distribution over 41 different domains. (b) Illustration of some video examples with question and answer labels on the “painting” and “salad” domains.
Table 1: Video question answering performances on different large video-language models. Following Video-ChatGPT [36], we employ ChatGPT-Assistant (gpt-3.5-turbo) to report the accuracy and score of the open-set answering.
Method LLM Size MSRVTT-QA ActivityNet-QA InternVid-QA
Acc Score Acc Score Acc Score
Video-ChatGPT 7B 49.3 2.8 35.2 2.7 - -
LLaMA-Adapter 7B 43.8 2.7 34.2 2.7 - -
LLaMA-VID 7B 57.7 3.2 47.4 3.3 - -
MiniGPT-4 7B - - - - 25.0 2.3
MiniGPT4-Video 7B 59.7 3.3 46.3 3.4 30.2 2.5
VideoChat2 7B 54.1 3.3 49.1 3.3 47.1 3.1
Video-LLaVA 7B 59.2 3.5 45.3 3.3 45.6 3.1
[1pt/1pt] Ours 7B 61.8 +2.6 3.5 46.9 +1.6 3.4 49.5 +3.9 3.1
4.2 Zero-shot Video Question Answering
We evaluate the advanced video language models on the open-ended video question-answering tasks under a zero-shot setting. The benchmark is composed of three datasets, involving ActivityNet-QA, MSRVTT-QA, and our newly proposed InternVid-QA. For a fair evaluation of the open-ended answers, we follow Video-ChatGPT [36] to use GPT-3.5 to evaluate the accuracy (True/False) and quality score (from 0 to 5). We compare the performances of several advanced large video-language models, including Video-ChatGPT [36], LLaMA-Adapter [61], Video-LLaVA [28], LLaMA-VID [27], MiniGPT-4 [65], MiniGPT4-Video [2] and VideoChat2 [25]. These competitors all conduct zero-shot video question answering conditioned on a single video. As mentioned above, our method’s answering is conditioned on the target video and several retrieved videos from the test set. According to the experimental results in Table 1, our framework improves the baseline (Video-LLaVA) by 2.6% on MSRVTT-QA, 1.6% on ActivityNet-QA and 3.9% on InternVid-QA, which shows its capability to effectively integrate multi-video knowledge and provide more reliable answers.
@@ 4.2 Zero-shot Video Question Answering
Table 2: Ablation study on video structuring and multi-video fusion components on InternVid-QA.
Struct Multi-video context $L$ Acc
single video 2.1K 45.6
multi-video tokens 12.5K 28.5
multi-video captions 2.5K 46.8
[1pt/1pt] ✓ single video 2.1K 46.1
graph fusion module 2.3K 49.5
Table 3: Ablation on the design of GFM.
HF-GAT CGA FFN Acc
46.2
47.8
49.5
48.7
@@ 4.3 Ablation study
Ablation study on the proposed components. Our framework mainly consists of two parts, which are video structuring and multi-video collaboration. We conduct ablation study on our proposed components to validate their effectiveness. Experimental results are shown in Table 2. We first disable video structuring module to test different common multi-video fusion strategies. Here “multi-video tokens” means concatenate all videos’ tokens as the input to context; “multi-video captions” means sending all videos’ captions to context. By observation, “multi-video tokens” strategy introduces overwhelming tokens to the context and brings excessive understanding challenges, yielding bad performance. “multi-video captions” can slightly improve the performance (1.2%) with the help of related videos. Then, we enable the Video Structuring Module and find that a structured and clear video representation is beneficial to LLM’s understanding, improving the performance by 0.5%. Finally, by adopting multi-video graph fusion, the framework can effectively extract useful information from related videos and get 3.4% additional performance improvement, at the cost of only 0.2K overhead of token numbers.
Ablation on Graph Fusion Module design. We further conduct an ablation study on the design of our Graph Fusion Module. Our GFM is composed of two components: HG-GAT and CGA. Since we have discussed the usage of FFN in Section 3.3, we also take it into consideration and conduct ablation study on these three components. Ablation results are shown in Table 3. In the first row, the graph structure feature tokens are directly sent to the multimodal projection layer to get the graph tokens. Then in the second row, we adopt HF-GAT to convey structure information to each token, increasing the performance by 1.6%. Then, with the help of CGA, multi-video knowledge is fused into graph tokens and further improves the performance by 1.7%. The adoption of FFN doesn’t bring improvement to the reasoning accuracy, which further corroborates the conclusion in Section 3.3 above.
Figure 6: Accuracy under different video numbers.
Figure 7: Visualization of our structured multi-video collaborative reasoning. We visualize a representative video question-answering example in our structured multi-video collaboration pipeline, including the grounding and structuring results, the cross-graph attention map, as well as the answers before and after adopting our framework.
Figure 8: Accuracy under different video relevance.
@@ 4.4 Discussion on the retrieved video contents
How do multiple videos affect the performance? Multi-video data can help to achieve more comprehensive reasoning, In this section, we discuss the effect of the number of the related videos. As illustrated in Figure 6, when the retrieval number increased from 0 to 7, we can observe that the accuracy first rises then drops, which peaks at 5.
How does video relevance affect the performance? In the main experiment, we arrange most related videos to the target video within each iteration. In this part, we rearrange the retrieved videos with different lower similarities (cosine similarity between video features) to test how video relevance affects the performance. As illustrated in Figure 8, the reasoning performance decreases as the relevance of the video diminishes, while the performance is still comparable to our competitors.
conclusion. More relevant videos can lead to better performance, but the key is to find videos containing critical information. Introducing more or unrelated videos will unavoidably introduce irrelevant information, yet our method possesses the ability to filter out irrelevant information to some extents.
@@ 4.5 Visualization
We visualize the reasoning process of our multi-video collaboration framework in Figure 7. The query is “What was the exact function of the machine used by the man?”, which requires high-level and domain-related knowledge and the baseline model failed to provide a satisfying response. In our framework, the video is represented as graph-structured data, maintaining the key spatio-temporal information. Then, through the cross-graph attention shown in Figure 7, important sub-graph structures from related videos are fused into current graph features, supporting our model to provide an accurate answer.

### v2

@@ 3.4 Structured Multi-Video Prompt
Upon obtaining the fused multi-video graph tokens, we need to integrate the graph, video, and text tokens together to create an LLM-friendly input. Therefore, we propose the structured multi-video prompt, as is shown in Fig. 5. Our prompt originates from previous video language models’ prompt design [lin2024video]. For the target video, we maintain the video tokens <<VIDEO_TOKENS>> of the target video to preserve fine-grained and detailed visual information. We also append its graph-based structured data <<GRAPH_TOKENS>> to indicate the key object and spatio-temporal relationships. For $N$ related videos, we only maintain the concise and data-efficient graph-based structured data <<GRAPH_TOKENS>>. In this context, we further indicate the relationships between the target video and related videos and how the LLMs can utilize these related multi-video structured data. By structuring the prompt in this manner, we enable the video language model to leverage multi-video information in an effective way, thereby enhancing the model’s ability to reason about and answer queries related to the video content.
@@ 4.1 Experiment Setup
Dataset descriptions. During the training phase, we construct a dataset containing structured video information for GFM training, based on the LLaVA-Video-178K dataset [zhang2024videoinstructiontuningsynthetic]. Specifically, we perform step-by-step preprocessing of the dataset as outlined in the Sec. 3.2. Notably, for video data in LLaVA-Video-178K that already includes captions, we retain the original captions without additional modifications to streamline the preprocessing workflow. Regarding video vectorization and retrieval mechanisms for related videos, we utilize Qwen3-Embedding-8B [zhang2025qwen3] to extract query embeddings (for retrieval) and document embeddings (for storage) from video captions. This approach is widely used in retrieval tasks [zhang2025qwen3]. Specifically, document embeddings are generated by directly inputting the captions, while query embeddings are generated with the designed prompt to prepare the input for the embedding model, as shown below:
Instruct: This is the caption of a video.
[…]
Query: {caption}.
Finally, we construct a training dataset comprising approximately 87K samples. While relatively small compared to datasets commonly used for training other VLMs (e.g., 87K vs. 9.36M [li2024llava]), our method achieves effective performance improvements, as shown in Tab. 2. This result highlights the ability of our method to integrate seamlessly into existing model frameworks and deliver performance gains through simple and efficient training on a compact dataset.
For evaluation, we test our approach on different video question-answering benchmarks, including ActivityNet-QA [yu2019activitynet], NExT-QA [xiao2021next], EgoSchema [mangalam2023egoschema], and Video-MME [fu2025video]. These benchmarks cover both short and long video understanding tasks, providing a comprehensive evaluation of our method.
Additionally, to enhance experimental efficiency for the ablation study, we utilize approximately 10% of the original training dataset for training. Similarly, during evaluation, we select subsets of the NExT-QA and EgoSchema datasets, each containing around 0.5K samples.
Implementation details. Our proposed structured multi-video collaboration framework is adaptive to general video language models. To verify the effectiveness of our proposed method, we conduct experiments with various models of different parameter scales on A6000 48GB GPUs, including LLaVA-OneVision-0.5B [li2024llava] and LLaVA-Video-7B [zhang2024videoinstructiontuningsynthetic]. For the Graph Fusion Module (GFM), the hidden state size is configured to match the output dimension of the corresponding vision encoder. We initialize the video language model (VLM) with pre-trained weights and further enhance its ability to effectively comprehend graph-based video representations by training it on our constructed dataset. To optimize our model efficiently, we adopt a standard two-stage training strategy [lin2024video, gao2024linvt]. During the first stage, we freeze the vision encoder, projector, and language model, focusing on training the GFM to align the inputs to the language model. In the second stage, we unfreeze the projector and language model while keeping the vision encoder frozen, and apply LoRA [hu2022lora] to the language model, then fine-tune the projector, GFM, and language model simultaneously. The detailed training recipe and hyperparameter configurations are provided in Tab. 1.
Table 1: The training recipe for VLM in our experiments.
Stage-1 Stage-2
Trainable GFM GFM, Projector, LLM
Batch size 128 64
Optimizer AdamW AdamW
Warmup ratio 0.03 0.03
Learning rate schedule Cosine decay Cosine decay
LR: $\boldsymbol{\phi}_{\text{GFM}}$ 1e-3 1e-4
LR: $\boldsymbol{\phi}_{\text{Proj.}}$ - 1e-5
LR: $\boldsymbol{\phi}_{\text{LLM}}$ - 1e-5
Table 2: Video question answering performances on different large video-language models. The wavy lines indicate the re-evaluated results.
Model Params Frames ActivityNet-QA NExT-QA EgoSchema Video-MME Average
Task Open-Ended Multi-Choice Multi-Choice Multi-Choice Acc. (%)
Duration Short Short Long Long
Video-LLaVA [lin2024video] 7B 8 45.30 62.60 38.40 40.40 46.68
LLaMA-VID [li2024llama] 7B 1fps 47.40 - 38.50 - -
PLLaVA [xu2024pllava] 7B 16 56.30 68.17 45.16 44.25 53.47
VideoChat2 [li2024mvbench] 7B 16 - - 54.40 47.90 -
LLaVA-NeXT-Video [zhang2024llavanextvideo] 7B 32 53.50 - 43.90 46.50 -
Qwen2-VL [wang2024qwen2] 7B 2fps 57.40 77.20 66.70 63.30 66.15
Qwen2.5-VL [bai2025qwen2] 3B 2fps - - 64.80 61.50 -
Qwen2.5-VL [bai2025qwen2] 7B 2fps - - 65.00 65.10 -
VideoLLaMA2 [cheng2024videollama] 7B 16 50.20 75.60 - 47.90 -
VideoLLaMA2.1 [cheng2024videollama] 7B 16 53.00 75.60 53.10 54.90 59.15
VideoLLaMA3 [zhang2025videollama] 2B 180 58.20 81.10 58.50 59.60 64.35
InternVL2 [chen2024expanding] 8B 16 - - 55.00 54.00 -
InternVL2.5 [chen2024expanding] 8B 64 58.90 85.00 51.50 64.20 64.90
NVILA [liu2025nvila] 8B 256 60.90 82.20 54.30 64.20 65.40
LLaVA-OneVision [li2024llava] 0.5B 32 45.65 57.20 26.80 44.00 43.41
   +Ours 0.5B 32 46.46 58.71 28.38 43.74 44.32
[1pt/1pt] LLaVA-Video [zhang2024videoinstructiontuningsynthetic] 7B 64 60.55 83.20 57.30 63.30 66.09
   +Ours 7B 64 61.25 84.00 61.76 64.37 67.84
4.2 Video Question Answering
We evaluate advanced video language models on video question-answering tasks, including ActivityNet-QA [yu2019activitynet], NExT-QA [xiao2021next], EgoSchema [mangalam2023egoschema], and Video-MME [fu2025video], which collectively span a diverse range of video understanding tasks. Following the methodology proposed by Video-LLaVA [lin2024video], we employ ChatGPT-Assistant to report the accuracy of open-ended answers for the ActivityNet-QA dataset. However, since the gpt-3.5-turbo-0613 model used in ActivityNet-QA’s original evaluation pipeline has been deprecated, for fair comparison, we opt to re-evaluate the results using Qwen3-235B-A22B [yang2025qwen3], an open-source large language model. Being open-source, this model is easier to access and utilize compared to the GPT series. Moreover, it exhibits superior language capabilities over the gpt-3.5-turbo series [yang2025qwen3], making it an ideal replacement for evaluating the correctness of open-ended answers. Thus, to ensure a fair and consistent comparison of baselines, we use Qwen3-235B-A22B to perform the re-evaluation of the ActivityNet-QA dataset for LLaVA-OneVision-0.5B and LLaVA-Video-7B. We further compare the performance of several advanced video language models, all of which perform video question answering conditioned on a single video. Unlike these traditional methods, our framework extends the reasoning capabilities by generating answers based on both the target video and multiple retrieved related videos. The experimental results, as presented in Tab. 2, highlight the superiority of our method over the baselines LLaVA-OneVision-0.5B and LLaVA-Video-7B. By introducing the concept of multi-video collaborative reasoning, our approach enhances average accuracy across various tasks, including open-ended answering, multiple-choice question answering, and video comprehension tasks covering videos of diverse durations. These results demonstrate that our method efficiently trains on the compact dataset, integrates multi-video knowledge, and provides more reliable answers.
@@ 4.2 Video Question Answering
Table 3: Ablation study on video structuring and multi-video fusion components on NExT-QA, conducted using the baseline model LLaVA-OneVision-0.5B [li2024llava].
Struct Multi-video context $L$ NExT-QA
single video 6.5K 61.4
multi-video tokens (32) 38K OOM
multi-video tokens (8) 15K 51.8
multi-video captions 9.3K 61.8
[1pt/1pt] ✓ single video 7.3K 62.0
graph fusion module 7.5K 65.2
Table 4: Ablation study on video structuring and multi-video fusion components on NExT-QA, conducted using the baseline model LLaVA-Video-7B [zhang2024videoinstructiontuningsynthetic].
Struct Multi-video context $L$ NExT-QA
single video 13K 79.8
multi-video tokens (64) 73K OOM
multi-video tokens (8) 22K 72.6
multi-video captions 16K 79.8
[1pt/1pt] ✓ single video 13.8K 83.6
graph fusion module 14K 84.2
Table 5: Ablation on the design of GFM. PA refers to Pooling Attention.
HF-GAT PA CGA FFN NExT-QA EgoSchema
61.4 26.4
[1pt/1pt] ✓ 64.2 28.0
64.4 28.2
65.0 28.6
64.4 27.6
Table 6: Ablations on different video retrieval strategies.
Video Retrieval Strategy NExT-QA EgoSchema
video vector-based retrieval 63.8 27.6
restricted retrieval 63.6 27.6
caption vector-based retrieval 65.0 28.6
(a) LLaVA-OneVision-0.5B.
(b) LLaVA-Video-7B.
Figure 6: Comparative analysis of accuracy (%) and context length (K) for NExT-QA across different models under varying numbers of related videos.
@@ 4.3 Ablation study
Ablation study on the proposed components. Our framework consists of two components: video structuring and multi-video collaboration. The experimental results are shown in Tab. 4 for LLaVA-OneVision-0.5B [li2024llava] and Tab. 4 for LLaVA-Video-7B [zhang2024videoinstructiontuningsynthetic]. First, we evaluate our framework by disabling the video structuring module and testing common multi-video fusion strategies. The “multi-video tokens” strategy involves concatenating all video tokens as input to the context, whereas the “multi-video captions” strategy sends the captions of all videos into the context. Initially, we use the default number of frames preset in the respective models for related videos, but this causes out-of-memory (OOM) issues during inference. This highlights that directly utilizing the “multi-video tokens” approach is impractical for real-world inference scenarios. We address this by reducing the number of frames for related videos to 8, while maintaining the default number of frames for the target video. Despite this adjustment, the “multi-video tokens” strategy introduces an excessive number of tokens into the context, challenging the model’s understanding and leading to noticeable performance degradation (-9.6% for LLaVA-OneVision-0.5B and -7.2% for LLaVA-Video-7B). Conversely, the “multi-video captions” approach demonstrates slight performance improvements (+0.4% for LLaVA-OneVision-0.5B), but performance remains unchanged for LLaVA-Video-7B. Enabling the Video Structuring Module provides a structured and clearer video representation, which helps large language models (LLMs) achieve better content understanding. This results in consistent performance improvements across models (+0.6% for LLaVA-OneVision-0.5B and +3.79% for LLaVA-Video-7B). Lastly, applying the multi-video graph fusion strategy allows the framework to extract valuable information from related videos effectively. This achieves substantial performance improvements (+3.8% for LLaVA-OneVision-0.5B and +4.4% for LLaVA-Video-7B) with minimal overhead, requiring only 0.2K additional tokens over single-video processing while leveraging the Video Structuring Module for enhanced reasoning.
Ablation on Graph Fusion Module design. We further conduct an ablation study on the design of our Graph Fusion Module, which comprises three components: HF-GAT, CGA, and the Pooling Attention in the TEM. To expedite both the training and evaluation, we perform these studies using the LLaVA-OneVision-0.5B [li2024llava] model, utilizing subsets of the original training and evaluation datasets to validate the effectiveness of the proposed components. Since we have discussed the usage of FFN in Sec. 3.3, we also take it into consideration and conduct the ablation study on these three components. The results of the ablation experiments are presented in Tab. 6. In the first row, the graph structure feature tokens are directly sent to the multimodal projection layer to get the graph tokens. In the second row, we incorporate HF-GAT to propagate structural information to each token, which increases performance by 2.8% on the NExT-QA dataset and 1.6% on the EgoSchema dataset. We also integrate Pooling Attention into TEM to embed structural information from specific scenes. The results demonstrate that incorporating this enriched information leads to a slight improvement compared to the original textual structural information (+3% on the NExT-QA, +1.8% on the EgoSchema). Furthermore, with the inclusion of CGA, multi-video knowledge is fused into graph tokens, resulting in further performance gains of 3.6% on the NExT-QA dataset and 2.2% on the EgoSchema dataset. Interestingly, the adoption of FFN does not improve reasoning accuracy, further supporting the conclusions drawn in Sec. 3.3.
(a) LLaVA-OneVision-0.5B.
(b) LLaVA-Video-7B.
Figure 7: Comparative analysis of accuracy (%) for NExT-QA across different models under varying relevance of related videos.
@@ 4.4 Discussion on the retrieved video contents
How do multiple videos affect the performance? Multi-video data contributes to achieving more comprehensive reasoning. In this section, we discuss the effect of the number of related videos. As shown in Fig. 6, increasing the number of retrieved videos from 1 to 8 causes accuracy to initially rise, peaking at 5 videos, before gradually declining. Importantly, this accuracy trend is accompanied by only a marginal increase in the total number of tokens.
How does video relevance affect the performance? In the above experiments, we arrange the most related videos to the target video within each iteration. In this part, we rearrange the retrieved videos with different lower similarities (measured as cosine similarity between retrieval features) to evaluate how video relevance impacts the performance. As illustrated in Fig. 7, the reasoning performance decreases as the relevance of the video diminishes, while the performance is still comparable to baseline.
How does the video retrieval strategy affect the performance? Intuitively, different retrieval strategies are the potential influencing factor for our collaborative reasoning method. Therefore, for a more detailed discussion, we conduct ablation studies on three video retrieval strategies:
  • Video vector-based retrieval. This strategy is applicable to most situations. It uses the SigLIP [zhai2023sigmoid] vision encoder to generate the features of the class tokens from sampled frames, then computes the average feature as each video’s feature vector, and constructs the video vector dataset. During inference, $N$ related videos are retrieved based on the highest cosine similarity between feature vectors.
  • Caption vector-based retrieval. This strategy applies to video datasets equipped with corresponding captions. It uses the text encoder to extract the feature vector from each video’s caption and constructs the caption vector dataset. During inference, $N$ related videos are retrieved by finding the highest cosine similarity between caption feature vectors.
  • Restricted retrieval. This strategy applies to artificially partitioned video datasets. Specifically, during inference, the retrieval process is restricted to videos within the test set, and the retrieval method follows the same procedure as caption vector-based retrieval.
We implement all three video retrieval strategies and test them on evaluation datasets. The results, presented in Tab. 6, indicate that the caption vector-based retrieval achieves the best performance, which can be attributed to the high-quality prompt construction (see Fig. 4) and the outstanding retrieval capabilities of Qwen3-Embedding [zhang2025qwen3]. Consequently, we adopt the caption vector-based retrieval strategy in this work. Nonetheless, the other strategies also demonstrate competitive performance, indicating that the reasoning process is only slightly affected by the choice of retrieval strategy. Overall, our framework exhibits robust performance across different video retrieval strategies.
Conclusion. Using more relevant videos can lead to better performance, as the key lies in retrieving videos that contain critical information. Incorporating more unrelated or less relevant videos inevitably introduces noise, but our method is capable of filtering out irrelevant information to some extent. Furthermore, our framework exhibits remarkable robustness, maintaining strong performance even when employing different video retrieval strategies.
Figure 8: Visualization of our structured multi-video collaborative reasoning. We present a representative video question-answering example from our structured multi-video collaboration pipeline, showcasing the Pooling Attention visualization, structuring results, and the Cross-Graph Attention map, along with the answers generated before and after applying our framework. Within each scene, color patches correspond to triplets with matching colors, highlighting regions of interest identified through Pooling Attention, while the dashed lines indicate the relationships between triplets within the scenes.
@@ 4.5 Visualization
We visualize the reasoning process of our multi-video collaboration framework in Fig. 8. The query is “What activities are the skateboarders performing in the video?”, which requires high-level and domain-related knowledge. The baseline model failed to provide a detailed response, offering only a generic description. In contrast, our framework represents the video as graph-structured data, preserving key spatio-temporal information. In Fig. 8, the color patches correspond to triplets with matching colors, highlighting key regions of interest identified through Pooling Attention. These regions emphasize the most significant visual features necessary for understanding the scene based on their associated triplets. Furthermore, the dashed lines illustrate the relationships between these triplets, demonstrating how structured video representations aggregate and pass based on the relational information across different frames and videos. Through the Cross-Graph Attention mechanism shown in Fig. 8, sub-graphs from related videos contribute useful relational structures to the current video’s graph. By fusing sub-graph features from related videos, our model builds a coherent understanding of complex scenarios, leading to an accurate and detailed response.
Figure 9: Visualization of video question answering examples.
4.6 More Video Question-Answering Results
We present additional video question-answering results in Fig. 9, which illustrate the strength of our framework in delivering accurate, detailed, and contextually precise answers, surpassing baseline methods.
For instance, our model effectively integrates domain knowledge to correctly interpret unique activities, as demonstrated by accurately identifying that the person in green clothes is “carving a watermelon to make a jack-o-lantern” rather than merely “cutting a watermelon with a knife.” Similarly, it provides more detailed and context-aware answers about the use of uneven bars by identifying them as “two parallel bars set at different heights” and specifying that the man is “using these bars to perform his routine.” Furthermore, our framework may alleviate hallucination problems to some extent, as shown in the case of evaluating safety in the bowling video, where it concludes that ”bowling is a safe sport as no injuries are shown” instead of generating fabricated details.
These examples demonstrate the robustness and reliability of our structured multi-video collaborative reasoning framework, which outperforms the baseline by providing precise, accurate, and context-aware answers across diverse queries.